OpenStudy (anonymous):

A high school basketball team has a budget specifically for towels and extra basketballs. A towel costs $3 and a basketball costs $18. Assume the budget is $450 and that 30 towels are needed for the team for the entire season. Use a standard form equation to determine how many basketballs may be purchased by the team. A. 5 B. 15 C. 20 D. 30

OpenStudy (anonymous):

Let the cost of the basketball be x Let the cost of the towel be y 18x + 3y = 450 the number of towels is given. y = 30 now find x

OpenStudy (anonymous):

could the answer be 15

OpenStudy (anonymous):

No. 15 is not the answer 18x + 3(30) = 450 18x + 90 = 450